# 2Pengcheng Laboratory

### v1

@@ Abstract
Deep learning-based point cloud modeling has been widely investigated as an indispensable component of general shape analysis. Recently, transformer and state space model (SSM) have shown promising capacities in point cloud learning. However, limited research has been conducted on medical point clouds, which have great potential in disease diagnosis and treatment. This paper presents an SSM-based hierarchical feature learning framework for medical point cloud understanding. Specifically, we down-sample input into multiple levels through the farthest point sampling. At each level, we perform a series of k-nearest neighbor (KNN) queries to aggregate multi-scale structural information. To assist SSM in processing point clouds, we introduce coordinate-order and inside-out scanning strategies for efficient serialization of irregular points. Point features are calculated progressively from short neighbor sequences and long point sequences through vanilla and group Point SSM blocks, to capture both local patterns and long-range dependencies. To evaluate the proposed method, we build a large-scale medical point cloud dataset named MedPointS for anatomy classification, completion, and segmentation. Extensive experiments conducted on MedPointS demonstrate that our method achieves superior performance across all tasks. The dataset is available at medpoints. Code is merged to a public medical imaging platform: flemme.
Keywords:
State Space Model Point cloud ModelingMedical Shape Analysis
**footnotetext: Guoqing Zhang and Jingyun Yang contributed equally to this work.
@@ 1 Introduction
Recently, the emergence of large-scale shape benchmarks [21, 1] has substantially accelerated the evolution of 3D vision. Qi et al. propose PointNet [14], which firstly introduces shared multilayer perceptron (MLP) and max pooling for global feature learning of point clouds. Subsequent works [15, 18] further improve performance by exploiting local context information. However, MLP and convolution have limited capabilities to capture long-range dependencies. Inspired by the immense success of transformer in natural language processing (NLP) [17, 2] and compute vision [3, 12], sequential modeling of point clouds has emerged as a dominant trend. Although transformer-based models [5, 24] have achieved remarkable progress in point cloud learning, the quadratic complexity of self-attention mechanisms restricts their scalability. On the other hand, one of the pioneers of state space models (SSM) named Mamba [4] has emerged as a popular network backbone in NLP and 2D vision tasks [11, 25, 19, 16] with linear complexity and powerful sequence representation ability. A few very recent studies [9, 10, 23] introduce grid and octree-based serialization to adapt Mamba for handling point clouds, achieving comparable performance to transformer-based counterparts while significantly reducing computational overhead.
Although there has been significant progress in general point cloud learning, few prior works systematically study medical point clouds due to the following challenges. First, the majority of medical data exists in the form of 2D or 3D images, resulting in a scarcity of large-scale medical point cloud datasets. The launch of MedShapeNet [8], a large-scale medical shape dataset, provides a promising foundation for addressing this problem. Second, medical point clouds inherently possess complex and nested structures, posing significant difficulties for the direct deployments of current methods.
@@ 2.2 SSM-based Point Feature Learning
[…]
@@
Here, $\operatorname{dist}(\mathcal{P},q)=\{\|p-q\|\,|\,p\in\mathcal{P}\}$, $\operatorname{mean}(\mathcal{P})$ is the mean coordinate of $\mathcal{P}$. As illustrated in Fig. 2 (a), inside-out scanning is preferable for capturing nested structures. Note that all scanning operations are performed in metric space. The transposed indices and inverted indices are denoted as $\Theta^{\top}$ and $\Theta^{-1}$, respectively. And we have:
@@ 3.1 MedPointS Dataset and Evaluation Metrics
MedShapeNet [8] is a large-scale medical shape dataset that contains over 100,000 meshes of various anatomical structures, from which we construct MedPointS, a large-scale dataset for medical point cloud understanding. Specifically, we select patients from MedShapeNet with relatively complete body part scans and compile a classification dataset comprising 28,737 anatomical structures across 46 categories, with each sample containing no more than 16,384 points. For the completion task, an anchor point is randomly selected from each point cloud, and 20% of the nearest points are removed to generate its incomplete counterpart. Furthermore, we integrate these anatomical structures based on their corresponding patients to create an anatomy segmentation dataset containing 1,020 samples, each of which has 65,536 points.
@@ 3.2 Experimental Setup
We compare our method with previous state-of-the-art point cloud learning models, including PointNet [14], PointNet++ [15], DGCNN [18], Point Cloud Transformer (PCT) [5] and vanilla PointMamba [9]. To ensure a fair comparison, we employ leaky ReLU activation and group normalization for all models. Each model contains 5 building blocks with feature channels of $[64,64,128,256,512]$ and one dense layer with an output channel $C_{z}=1024$. Point clouds are randomly sampled to a size of 2,048, 2,048, and 4,096 for classification, completion, and segmentation tasks, respectively. Classification models are trained using Adam [7] optimizer for 100 epochs, while completion and segmentation models are trained for 200 epochs. All models are trained on an NVIDIA A800 GPU. The learning rate starts from $1\text{e}^{-4}$ and follows a linear decaying schedule.
@@ 3.3 Results
Table 2: Ablation studies. $\mathbf{S}$ and $\mathbf{D}$ indicate static ball query in metric space and dynamic KNN query in feature space, respectively. The best results are denoted in Bold. Default settings are marked with grey.
@@ 4 Conclusion
In this paper, we first present a comprehensive study on medical point cloud understanding. We propose an SSM-based learning framework with a hierarchical design to integrate multi-scale neighbor information. We introduce coordinate-order and inside-out scanning strategies to serialize points into short neighbor sequences and long sequences, which are further processed by vanilla and group PSSM blocks to capture both local and global contextual relationships. In addition, we compile a large-scale medical point cloud dataset named MedPointS, to benchmark model performance on anatomy classification, completion, and segmentation tasks. Extensive experiments conducted on MedPointS showcase the reliable capabilities of our method in modeling medical point clouds.
Although our method achieves notable progress in medical point cloud learning, several aspects remain worthy of further exploration. Current models are trained exclusively on point clouds, while integrating image and language contexts is a promising direction for future enhancements. Moreover, the effectiveness of SSMs in handling complex tubular structures with distinct topological features is still unclear. In addition, the potential of transformers and SSMs in advancing medical shape generation has yet to be thoroughly investigated. These areas will be the primary focus of our future research efforts.

### v2

# Hierarchical Feature Learning for Medical Point Clouds via State Space Model

@@ Abstract
Deep learning-based point cloud modeling has been widely investigated as an indispensable component of general shape analysis. Recently, transformer and state space model (SSM) have shown promising capacities in point cloud learning. However, limited research has been conducted on medical point clouds, which have great potential in disease diagnosis and treatment. This paper presents an SSM-based hierarchical feature learning framework for medical point cloud understanding. Specifically, we down-sample the input into multiple levels through the farthest point sampling. At each level, we perform a series of k-nearest neighbor (KNN) queries to aggregate multi-scale structural information. To assist SSM in processing point clouds, we introduce coordinate-order and inside-out scanning strategies for efficient serialization of irregular points. Point features are calculated progressively from short neighbor sequences and long point sequences through vanilla and group Point SSM blocks, to capture both local patterns and long-range dependencies. To evaluate the proposed method, we build a large-scale medical point cloud dataset named MedPointS for anatomy classification, completion, and segmentation. Extensive experiments conducted on MedPointS demonstrate that our method achieves superior performance across all tasks. The dataset is available at https://flemme-docs.readthedocs.io/en/latest/medpoints.html. Code is merged into a public medical imaging platform: https://github.com/wlsdzyzl/flemme.
@@ 1 Introduction
Recently, the emergence of large-scale shape benchmarks [21, 1] has substantially accelerated the evolution of 3D vision. Qi et al. propose PointNet [14], which firstly introduces shared multilayer perceptron (MLP) and max pooling for global feature learning of point clouds. Subsequent works [15, 18] further improve performance by exploiting local context information. However, MLP and convolution have limited capabilities to capture long-range dependencies. Inspired by the immense success of transformers in natural language processing (NLP) [17, 2] and computer vision [3, 12], sequential modeling of point clouds has emerged as a dominant trend. Although transformer-based models [5, 25] have achieved remarkable progress in point cloud learning, the quadratic complexity of self-attention mechanisms restricts their scalability. On the other hand, one of the pioneers of state space models (SSM) named Mamba [4] has emerged as a popular network backbone in NLP and 2D vision tasks [11, 26, 19, 16] with linear complexity and powerful sequence representation ability. A few very recent studies [9, 10, 24] introduce grid and octree-based serialization to adapt Mamba for handling point clouds, achieving comparable performance to transformer-based counterparts while significantly reducing computational overhead.
Although there has been significant progress in general point cloud learning, few prior works systematically study medical point clouds due to the following challenges. First, the majority of medical data exists in the form of 2D or 3D images, resulting in a scarcity of large-scale medical point cloud datasets. The launch of MedShapeNet [8], a large-scale medical shape dataset, provides a promising foundation for addressing this problem. Second, medical point clouds inherently possess complex and nested structures, posing significant difficulties for the direct deployment of current methods.
@@ 2.2 SSM-based Point Feature Learning
$\Theta_{\text{z}}=\arg\operatorname{sort}(\mathcal{P}[\text{'z'}]),$ (3)
where $\mathcal{P}[\text{'z'}]$ returns the z-coordinates of point set $\mathcal{P}$, $\Theta_{\text{z}}$ is the sorted indices. Similarly, we have x-order scanning $\Theta_{\text{x}}$ and y-order scanning $\Theta_{\text{y}}$. Considering that medical point clouds often contain multiple nested structures, we introduce a simple inside-out scanning strategy that can be formulated as:
@@
Here, $\operatorname{dist}(\mathcal{P},q)=\{\|p-q\|\,|\,p\in\mathcal{P}\}$, $\operatorname{mean}(\mathcal{P})$ is the mean coordinate of $\mathcal{P}$. As illustrated in Fig. 2 (a), inside-out scanning is preferable for capturing nested structures. Note that all scanning operations are performed in metric space. The transposed indices and inverted indices are denoted as $\Theta^{\top}$ and $\Theta^{-1}$, respectively, with the property:
@@ 3.1 MedPointS Dataset and Evaluation Metrics
MedShapeNet [8] is a large-scale medical shape dataset that contains over 100,000 meshes of various anatomical structures, from which we construct MedPointS, a large-scale dataset for medical point cloud understanding. Specifically, we selected patients from MedShapeNet who had relatively complete body part scans, with the requirement of full vertebral coverage. We then constructed a classification dataset comprising 28,737 anatomical structures across 46 categories, where each sample contained no more than 16,384 points. Cases with corrupted or incomplete files were systematically excluded during the processing. For the completion task, an anchor point is randomly selected from each point cloud, and 20% of the nearest points are removed to generate its incomplete counterpart. Furthermore, we integrate these anatomical structures based on their corresponding patients to create an anatomy segmentation dataset containing 1,020 samples, each of which has 65,536 points.
@@ 3.2 Experimental Setup
We compare our method with previous state-of-the-art point cloud learning models, including PointNet [14], PointNet++ [15], DGCNN [18], Point Cloud Transformer (PCT) [5] and vanilla PointMamba [9]. To ensure a fair comparison, we employ leaky ReLU activation and group normalization for all models. Each model contains 5 building blocks with feature channels of $[64,64,128,256,512]$ and one dense layer with an output channel $C_{z}=1024$. Point clouds are randomly sampled to a size of 2,048, 2,048, and 4,096 for classification, completion, and segmentation tasks, respectively. Classification models are trained using Adam [7] optimizer for 100 epochs, while completion and segmentation models are trained for 200 epochs. All models are trained on an NVIDIA A800 GPU. The learning rate starts from $1\text{e}^{-4}$ and follows a linear decaying schedule. All experiments are conducted based on a public medical imaging platform Flemme [23].
@@ 3.3 Results
Table 2: Ablation studies. $\mathbf{S}$ and $\mathbf{D}$ indicate static ball query in metric space and dynamic KNN query in feature space, respectively. The best results are denoted in Bold. Default settings are marked in grey.
@@ 4 Conclusion
In this paper, we first present a comprehensive study on medical point cloud understanding. We propose an SSM-based learning framework with a hierarchical design to integrate multi-scale neighbor information in dynamic feature spaces. We introduce coordinate-order and inside-out scanning strategies to serialize points into short neighbor sequences and long sequences, which are further processed by vanilla and group PSSM blocks to capture both local and global contextual relationships. In addition, we compile a large-scale medical point cloud dataset named MedPointS, to benchmark model performance on anatomy classification, completion, and segmentation tasks. Extensive experiments conducted on MedPointS showcase the reliable capabilities of our method in modeling medical point clouds.
{credits}
4.0.1 Acknowledgements
This work is supported in part by the Natural Science Foundation of China (Grant 62371270).
[…]
The authors have no conflicts of interest to declare that are relevant to the content of this article.